# biblio.bib

### v1

@@ Abstract
Efficiently compiling quantum operations remains a major bottleneck in scaling quantum computing. Today’s state-of-the-art methods achieve low compilation error by combining search algorithms with gradient-based parameter optimization, but they incur long runtimes and require multiple calls to quantum hardware or expensive classical simulations, making their scaling prohibitive. Recently, machine-learning models have emerged as an alternative, though they are currently restricted to discrete gate sets. Here, we introduce a multimodal denoising diffusion model that simultaneously generates a circuit’s structure and its continuous parameters for compiling a target unitary. It leverages two independent diffusion processes, one for discrete gate selection and one for parameter prediction. We benchmark the model over different experiments, analyzing the method’s accuracy across varying qubit counts, circuit depths, and proportions of parameterized gates. Finally, by exploiting its rapid circuit generation, we create large datasets of circuits for particular operations and use these to extract valuable heuristics that can help us discover new insights into quantum circuit synthesis. Code, models and tutorials are publicly available at https://github.com/FlorianFuerrutter/genQC.
1 Introduction
@@
Synthesizing quantum circuits for given quantum operations is a highly non-trivial task, and stands as one of the key challenges in the pursuit of large-scale quantum computation. As quantum hardware continues to improve, with increasing qubit counts and lower error rates, we move closer to regimes where quantum advantage may become feasible. Indeed, quantum computational advantage has now been widely demonstrated, as for instance in Shor’s factoring algorithm [shor1994algorithms] and Grover’s search [grover1996fast], or in applications such as optimization [munoz2025scaling] and machine learning [liu2021rigorous]. However, many of these algorithms rely on large, fault-tolerant quantum computers, which remain out of reach. Despite notable advances in hardware, we are still in the era of Noisy Intermediate-Scale Quantum (NISQ) devices [preskill2018quantum], where limited qubit connectivity and different error sources hinder quantum computation. Another practical challenge arises from the diversity of quantum computing paradigms, from gate-based quantum computers realized by photonic [madsen2022quantum], superconducting [arute2019quantum], neutral atoms [bluvstein2024logical] and trapped ions [monz2016realization] architectures to measurement-based quantum computation [briegel2009measurement]. While this variety is promising and enriches the field, it also introduces difficulties: each platform features different native gate sets and qubit connectivity constraints, meaning that the optimal compilation of a given operation can vary significantly from one architecture to another.
@@
To address these challenges, various powerful compilation techniques have been developed, typically involving complex pipelines composed of multiple modules [younis2021qfast, madden2022best, grimsley2019adaptive]. While they are able to output highly accurate circuits, these methods tend to be slow and rely on heuristics, search algorithms and gradient-based optimization, often resulting in runtimes that scale prohibitively with qubit count [davis2020towards, madden2022best]. Nonetheless, from a fundamental and theoretical perspective, improving circuit synthesis depends on a deeper understanding of quantum circuits and how different gate combinations give rise to different computations. Beyond a purely theoretical study, achieving this requires methods that can generate circuits not only accurately but also efficiently, enabling the creation of datasets from which, when combined with expert knowledge in quantum information and computation, better compilation strategies can be discovered.
@@
    We present a method that compiles input unitaries into quantum circuits, predicting simultaneously the circuit’s structure (i.e., gate types and distribution) and its continuous parameters. It is based on a multimodal diffusion model that couples two independent diffusion processes: one addresses the discrete task of selecting gate types, while the other proposes the corresponding continuous gate parameters.
@@
    We benchmark our model on the unitary compilation problem, showcasing its capabilities across varying qubit counts and circuit depths.
@@
    Leveraging the fast generation capabilities of our method, we generate large circuit datasets for specific unitaries and use these to reveal interesting structural patterns.
2 Preliminaries
@@
The unitary compilation problem is defined by two inputs: a unitary matrix $\mathcal{U}$ of size $2^{n}\times 2^{n}$, where $n$ is the number of qubits, and a finite gate set, often related to the available physical operations of the quantum hardware onto which we aim to compile the unitary. Gates can be either discrete (e.g., the Hadamard or CNOT gates) or parameterized, in which case they are defined by a type and a parameter $\theta$ that can vary continuously (e.g., the X-rotation gate $R_{x}(\theta)$, with $\theta\in[0,4\pi]$ [nielsen2010quantum]). The output of the problem is a quantum circuit, i.e., a sequence of quantum gates drawn from the specified gate set that implements $\mathcal{U}$ on quantum hardware. Given a circuit $C$, we define its distance to the target unitary $\mathcal{U}$ via its unitary representation $\mathcal{U}_{C}$, using the infidelity:
$\mathcal{I}=1-\frac{1}{4^{n}}|\mathrm{Tr}(\mathcal{U}^{\dagger}_{C}\mathcal{U})|^{2},\text{ with }\mathcal{I}=0\text{ for }\mathcal{U}_{C}=\mathcal{U}\text{up to a global phase, and }\ 0\leq\mathcal{I}\leq 1.$ (1)
@@
Unitary compilation is a fundamental task in quantum computation and has therefore received significant attention. While the Solovay-Kitaev theorem ensures efficient compilation for single-qubit gates [dawson2005solovay], no guarantees exist for multi-qubit systems. In this regime, a variety of heuristic, search-based, gradient-based, and machine learning methods have been explored. Approaches such as [davis2020towards, madden2022best, du2022quantum, grimsley2019adaptive] alternate between searching for a circuit structure and optimizing gate parameters until they reach a desired error threshold. These methods achieve high fidelity but at the expense of long runtimes and deep circuits. In recent years, machine learning has had a substantial impact on the field [krenn2023artificial]. For example, Ref. [he2023gnn] proposed the use of graph neural networks to improve the ansatz selection within the aforementioned iterative framework, while [ostaszewski2021reinforcement] achieves this via reinforcement learning. In the latter, an agent is trained to sequentially place gates in a circuit and is rewarded when the resulting unitary closely approximates the target (see also  [bolens2021reinforcement, melnikov2018active, rietsch2024unitary, moro2021quantum, olle2024simultaneous, preti2024hybrid]).
@@
Importantly, all previous methods consider cost functions or rewards that require the simulation of the proposed circuit, a classically hard computation which motivated the development of methods working natively on quantum computers [khatri2019quantum]. Another approach to circumvent this bottleneck is to train a generative model on a dataset of labeled circuits, as done for instance in Ref. [nakaji2024generative]. Although generating and labeling such a dataset is computationally expensive, once it is built, all further training is purely classical. In this direction, Ref. [furrutter2024quantum] proposed the use of DMs for quantum circuit synthesis, albeit demonstrated only on discrete gate sets and circuits with up to three qubits for unitary compilation.
@@
Diffusion models (DM) are a class of generative models that aim to learn the underlying probability distribution $p_{\text{data}}(\mathbf{x})$ of a dataset. Starting from a clean data sample $\mathbf{x}$, the forward diffusion process defines a sequence of latent variables $\{\mathbf{z}_{t}\}_{t\in[0,1]}\coloneq\mathbf{z}_{0:1}$, where each $\mathbf{z}_{t}$ is a progressively noisier version of $\mathbf{x}$ starting at $t=0$. The objective of the model is to recover the original sample from a noisy observation $\mathbf{z}_{t}$. In continuous domains, this process can be formalized either as a stochastic differential equation (SDE) [karras2022elucidating] or equivalently as a discrete-time Markov chain [ho2020denoising]. In the case of Gaussian diffusion, the transition probability from $\mathbf{z}_{t}$ given $\mathbf{x}$ is defined as
@@
Diffusion models are trained by maximizing the evidence lower bound (ELBO) of the log-likelihood of $\mathbf{x}$  [ho2020denoising]. Once the DM is trained, sampling is typically performed via ancestral sampling: starting from a fully complete noisy latent $\mathbf{z}_{1}\sim\mathcal{N}(\mathbf{0},\mathbf{I})$, one follows the reverse transitions $p(\mathbf{z}_{s}|\mathbf{z}_{t})$ for $0\leq s<t\leq 1$ until reaching a clean sample at $t=0$, as for instance described in Ref. [ho2020denoising].
@@
Multimodal diffusion models (DMs) address the generation of samples composed of multiple modes. Recent works explored embedding multimodal data into a joint latent representation and applying a single Gaussian diffusion process [chen2024diffusion, si2025tabrep]. Other approaches rather consider two identical Gaussian processes [bao2023one, ruan2023mm]. On the other hand, previous works have also explored generation via separated diffusion processes as, e.g., in Ref. [kotelnikov2023tabddpm] by combining Gaussian with multinomial diffusion or in Ref. [shi2024tabdiff] using a Gaussian and a masked diffusion process.
3 Methods
Figure 1: Multimodal quantum circuit synthesis pipeline scheme. a) An input circuit is first tokenized and then embedded into two separate modes, from which the forward and backward diffusion processes are defined. b) Schematic of the generative model. See Section 3.1 and Section 3.2 for details.
3.1 Multimodal diffusion process
To perform simultaneous generation of discrete and parameterized gates, we represent each quantum circuit $\mathbf{x}$ as a combination of two different modes: a discrete (categorical) mode $\mathbf{h}_{0}$ that encodes the gate types, and a continuous mode $\mathbf{w}_{0}$ that specifies the values of the parameterized gates. We then design the forward process of the DM as two independent diffusion processes, each acting on its respective data mode (Fig. 1a). This separation allows us to construct mode-specific embeddings, tailored to the characteristics of each data type. In particular, we set the joint probability of the trajectories $\mathbf{h}_{0:1}$ and $\mathbf{w}_{0:1}$ of the forward process as the product of the independent joint distributions:
@@ 3.1 Multimodal diffusion process
where the separate distributions are defined via the standard Gaussian diffusion (see Eq. 2). Although the forward process in Eq. 2 is already component-independent, we further consider here that each diffusion process has a different noise schedule, namely $\bar{\alpha}_{t}^{h}$ and $\bar{\alpha}_{t}^{w}$, with the corresponding signal-to-noise ratios $\mathrm{SNR}_{h}(t)\coloneq\bar{\alpha}_{t}^{h}/(1-\bar{\alpha}_{t}^{h})$ and $\mathrm{SNR}_{w}(t)\coloneq\bar{\alpha}_{t}^{w}/(1-\bar{\alpha}_{t}^{w})$.

### v2

# Synthesis of discrete-continuous quantum circuits with multimodal diffusion models

@@ Abstract
Efficiently compiling quantum operations remains a major bottleneck in scaling quantum computing. Today’s state-of-the-art methods achieve low compilation error by combining search algorithms with gradient-based parameter optimization, but they incur long runtimes and require multiple calls to quantum hardware or expensive classical simulations, making their scaling prohibitive. Recently, machine-learning models have emerged as an alternative, though they are currently restricted to discrete gate sets. Here, we introduce a multimodal denoising diffusion model that simultaneously generates a circuit’s structure and its continuous parameters for compiling a target unitary. It leverages two independent diffusion processes, one for discrete gate selection and one for parameter prediction. We benchmark the model over different experiments, analyzing the method’s accuracy across varying qubit counts and circuit depths, showcasing the ability of the method to outperform existing approaches in gate counts and under noisy conditions. Additionally, we show that a simple post-optimization scheme allows us to significantly improve the generated ansätze. Finally, by exploiting its rapid circuit generation, we create large datasets of circuits for particular operations and use these to extract valuable heuristics that can help us discover new insights into quantum circuit synthesis. Code and models are publicly available at https://github.com/FlorianFuerrutter/genQC.
I Introduction
@@
Synthesizing quantum circuits for given quantum operations is a highly non-trivial task, and stands as one of the key challenges in the pursuit of large-scale quantum computation. As quantum hardware continues to improve, with increasing qubit counts and lower error rates, we move closer to regimes where quantum advantage may become feasible. Indeed, quantum computational advantage has now been widely demonstrated, as for instance in Shor’s factoring algorithm [1] and Grover’s search [2], or in applications such as optimization [3] and machine learning [4]. However, many of these algorithms rely on large, fault-tolerant quantum computers, which remain out of reach. Despite notable advances in hardware, we are still in the era of Noisy Intermediate-Scale Quantum (NISQ) devices [5], where limited qubit connectivity and different error sources hinder quantum computation. Another practical challenge arises from the diversity of quantum computing paradigms, from gate-based quantum computers realized by photonic [6], superconducting [7], neutral atoms [8] and trapped ions [9] architectures to measurement-based quantum computation [10]. While this variety is promising and enriches the field, it also introduces difficulties: each platform features different native gate sets and qubit connectivity constraints, meaning that the optimal compilation of a given operation can vary significantly from one architecture to another.
@@
To address these challenges, various powerful compilation techniques have been developed, typically involving complex pipelines composed of multiple modules [11, 12, 13, 14]. While they are able to output highly accurate circuits, these methods tend to be slow and rely on heuristics, search algorithms and gradient-based optimization, often resulting in runtimes that scale prohibitively with qubit count [15, 13]. Nonetheless, from a fundamental and theoretical perspective, improving circuit synthesis depends on a deeper understanding of quantum circuits and how different gate combinations give rise to different computations. Beyond a purely theoretical study, achieving this requires methods that can generate circuits not only accurately but also efficiently, enabling the creation of datasets from which, when combined with expert knowledge in quantum information and computation, better compilation strategies can be discovered.
@@
    We present a method that synthesizes input unitaries into quantum circuits, predicting simultaneously the circuit’s structure (i.e., gate types and distribution) and its continuous parameters. It is based on a multimodal diffusion model that couples two independent diffusion processes: one addresses the discrete task of selecting gate types, while the other proposes the corresponding continuous gate parameters.
@@
    We benchmark our model on the unitary compilation problem, showcasing its capabilities across varying qubit counts and circuit depths. When comparing to existing approaches, our method excels at generating shorter circuits, although at the expense of lower accuracies, making it a promising tool for NISQ devices. We then present a simple search approach that boosts the accuracy of the method.
@@
    Leveraging the fast generation capabilities of our method, we generate large circuit datasets for specific unitaries and use these to reveal interesting structural patterns. For instance, our method is able to find the textbook circuit for the quantum Fourier transform.
II Preliminaries
@@
The unitary compilation problem is defined by two inputs: a complex-valued unitary matrix $\mathcal{U}$ of size $2^{n}\times 2^{n}$, where $n$ is the number of qubits, and a finite gate set, often related to the available physical operations of the quantum hardware onto which we aim to compile the unitary. Gates can be either fixed (e.g., the Hadamard or CNOT gates) or parameterized, in which case they are defined by a type and a parameter $\theta$ that can vary continuously (e.g., the X-rotation gate $R_{x}(\theta)$, with $\theta\in[0,4\pi]$ [16]). In the following, we will also refer to the former as discrete gates, and to the latter as both parametrized and continuous gates. We note that any discrete gate can be decomposed into continuous gates, but allowing the usage of both gate types can help get better and more interpretable circuits (see Fig.˜8). The output of the compilation problem is a quantum circuit, i.e., a sequence of quantum gates drawn from the specified gate set that implements $\mathcal{U}$ on quantum hardware. Given a circuit $C$, we define its distance to the target unitary $\mathcal{U}$ via its unitary representation $\mathcal{U}_{C}$, using the infidelity:
$\mathcal{I}=1-\frac{1}{4^{n}}|\mathrm{Tr}(\mathcal{U}^{\dagger}_{C}\mathcal{U})|^{2},$ (1)
@@
with $\mathcal{I}=0$ for $\mathcal{U}_{C}=\mathcal{U}$ up to a global phase, and $0\leq\mathcal{I}\leq 1$. Unitary compilation is a fundamental task in quantum computation and has therefore received significant attention. While the Solovay-Kitaev theorem ensures efficient compilation for single-qubit gates [17], no guarantees exist for multi-qubit systems. In this regime, a variety of heuristic, search-based, gradient-based, and machine learning methods have been explored. Approaches such as [15, 13, 18, 14] alternate between searching for a circuit structure and optimizing gate parameters until they reach a desired error threshold. These methods achieve high fidelity but at the expense of long runtimes and deep circuits. In recent years, machine learning has had a substantial impact on the field [19]. For example, Ref. [20] proposed the use of graph neural networks to improve the ansatz selection within the aforementioned iterative framework, while [21] achieves this via reinforcement learning (RL). In the latter, an agent is trained to sequentially place gates in a circuit and is rewarded when the resulting unitary closely approximates the target (see also  [22, 23, 24, 25, 26, 27]).
@@
Importantly, all previous methods consider cost functions or rewards that require the simulation of the proposed circuit, a classically hard computation which motivated the development of methods working natively on quantum computers [28]. Another approach to circumvent this bottleneck is to train a generative model on a dataset of labeled circuits, as done for instance in Ref. [29]. Although generating and labeling such a dataset is computationally expensive, once it is built, all further training is purely classical. In this direction, Ref. [30] proposed the use of DMs for quantum circuit synthesis, albeit demonstrated only on discrete gate sets and circuits with up to three qubits for unitary compilation.
@@
Diffusion models (DM) are a class of generative models that aim to learn the underlying probability distribution $p_{\text{data}}(\mathbf{x})$ of a dataset. Starting from a clean data sample $\mathbf{x}$, the forward diffusion process defines a sequence of latent variables $\{\mathbf{z}_{t}\}_{t\in[0,1]}\coloneq\mathbf{z}_{0:1}$, where each $\mathbf{z}_{t}$ is a progressively noisier version of $\mathbf{x}$ starting at $t=0$. The objective of the model is to recover the original sample from a noisy observation $\mathbf{z}_{t}$. In continuous domains, this process can be formalized either as a stochastic differential equation (SDE) [31] or equivalently as a discrete-time Markov chain [32]. In the case of Gaussian diffusion, the transition probability from $\mathbf{z}_{t}$ given $\mathbf{x}$ is defined as
@@
Diffusion models are trained by maximizing the evidence lower bound (ELBO) of the log-likelihood of $\mathbf{x}$  [32]. Once the DM is trained, sampling is typically performed via ancestral sampling: starting from a fully complete noisy latent $\mathbf{z}_{1}\sim\mathcal{N}(\mathbf{0},\mathbf{I})$, one follows the reverse transitions $p(\mathbf{z}_{s}|\mathbf{z}_{t})$ for $0\leq s<t\leq 1$ until reaching a clean sample at $t=0$, as for instance described in Ref. [32].
Figure 1: Multimodal quantum circuit synthesis pipeline scheme. a) An input circuit is first tokenized and then embedded into two separate modes, from which the forward and backward diffusion processes are defined. b) Schematic of the generative model. c) Inference overview, exemplary for the quantum Fourier transform. See Section˜III for details.
@@
Multimodal diffusion models (DMs) address the generation of samples composed of multiple modes. Recent works explored embedding multimodal data into a joint latent representation and applying a single Gaussian diffusion process [33, 34]. Other approaches rather consider two identical Gaussian processes [35, 36]. On the other hand, previous works have also explored generation via separated diffusion processes as, e.g., in Ref. [37] by combining Gaussian with multinomial diffusion or in Ref. [38] using a Gaussian and a masked diffusion process.
III Methods
III.1 Multimodal diffusion process
To perform simultaneous generation of discrete and parameterized gates, we represent each quantum circuit $\mathbf{x}$ as a combination of two different modes: a discrete (categorical) mode $\mathbf{h}_{0}$ that encodes the gate types, and a continuous mode $\mathbf{w}_{0}$ that specifies the values of the parameterized gates. We then design the forward process of the DM as two independent diffusion processes, each acting on its respective data mode (Fig.˜1a). This separation allows us to construct mode-specific embeddings, tailored to the characteristics of each data type. In particular, we set the joint probability of the trajectories $\mathbf{h}_{0:1}$ and $\mathbf{w}_{0:1}$ of the forward process as the product of the independent joint distributions:
@@ III.1 Multimodal diffusion process
where the separate distributions are defined via the standard Gaussian diffusion (see Eq.˜2). Although the forward process in Eq.˜2 is already component-independent, we further consider here that each diffusion process has a different noise schedule, namely $\bar{\alpha}_{t}^{h}$ and $\bar{\alpha}_{t}^{w}$, with the corresponding signal-to-noise ratios $\mathrm{SNR}_{h}(t)\coloneq\bar{\alpha}_{t}^{h}/(1-\bar{\alpha}_{t}^{h})$ and $\mathrm{SNR}_{w}(t)\coloneq\bar{\alpha}_{t}^{w}/(1-\bar{\alpha}_{t}^{w})$.